The length of a rectangle is 6cm longer than its width. If the perimeter of the rectangle is 32cm, find its area. How Would you do this?
Answer by rapaljer(4671)

Let x = width of the rectangle
x+6 = length of the rectangle

Perimeter = 2(W) + 2(L) = 32
2(x) + 2(x+6) = 32
2x + 2x + 12 = 32
4x + 12 = 32
4x = 20
x= 5 cm Width
x+6 = 11 cm Length


Area = L*W = 11 cm * 5 cm = 55 sq cm.

R^2 at SCC